OpenStudy (anonymous):

Solve x+12/x+4=x/x+8

OpenStudy (lgbasallote):

\(\LARGE \frac{x+12}{x+4} = \frac{x}{x+8}\) that's the question right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

why don't you try cross multiplying..you know how right?

OpenStudy (anonymous):

Yeah i did that but I keep getting stuck. i got down to x^2+20x+20=x^2+4x and then just got stuck

OpenStudy (lgbasallote):

good one! now put x^2 + 4x on to the left side....subtract x^2 from both sides first what do you have?

OpenStudy (anonymous):

oh okay i tried that. so20x+20=4x oh wait then you have to subtract 20x from both sides rights?

OpenStudy (lgbasallote):

no no..you always subtract the SMALLER one...so subtract 4x from both sides

OpenStudy (anonymous):

yeah but if you do that then there is nothing left on the left side.. how does that work?

OpenStudy (lgbasallote):

btw..there was a problem with your (x+12)(x+8) that should be equal to x^2 + 20x + 96 because 12 x 8 is 96 NOT 20..got it? :)

OpenStudy (anonymous):

ohh i added it oops :P okay lemme start over

OpenStudy (lgbasallote):

no no..it's okay..let's carry on from 20x + 96 = 4x

OpenStudy (lgbasallote):

there wasnt any big changes on the 96 anyway..so it wont influence our solution so far

OpenStudy (lgbasallote):

so...you subtract both sides by 4x what do you have?

OpenStudy (anonymous):

16x+96 but i didnt think you could leave one side with nothing

OpenStudy (lgbasallote):

where'd that equal sign go? what's 4x - 4x?

OpenStudy (anonymous):

... 0 ha okay i feel kinda dumb so 16x+96=0 then you subtract 96?

OpenStudy (lgbasallote):

RIGHT!!! \(\checkmark\) :DDD

OpenStudy (anonymous):

and then 16/-96=-6?

OpenStudy (lgbasallote):

you got it! :DDD

OpenStudy (lgbasallote):

but that's -96/16 but i think you got the idea haha

OpenStudy (anonymous):

Thanks :) oh ha yeah